# Understanding Flow Behaviors of Supercooled Liquids by Embodying Solid-Liquid Duality at Particle Level

### v1

@@ Abstract
Supercooled liquids exhibit intricate flow behaviors, which progressively become nonlinear as flow rate increases. Conceptually, this complexity can be understood by the solid-liquid duality in Maxwell’s understanding of materials’ response to external load. Nevertheless, the microscopic foundation of this duality in supercooled liquids remains elusive, thereby impeding the modeling of flow behaviors from a microscopic viewpoint. The existence of dynamic heterogeneity adds to this difficulty. To tackle these problems, we propose the concept of local configurational relaxation time $\tau_{\mathrm{LC}}$, which is defined at the particle level. The spatial distribution of $\tau_{\mathrm{LC}}$ in flow is heterogeneous. Depending on the comparison between the local mobility measured by $\tau_{\mathrm{LC}}$ and the external shear rate, the shear response of local regions is either solid-like or liquid-like. In this way, $\tau_{\mathrm{LC}}$ plays a role similar to the Maxwell time. By applying this microscopic solid-liquid duality to different conditions of shear flow with a wide range of shear rates, we describe the emergence of shear thinning in steady shear, and predict the major characteristics of the transient response to start-up shear. Furthermore, we reveal a clear structural foundation for $\tau_{\mathrm{LC}}$ and the solid-liquid duality associated with it by introducing an order parameter extracted from local configuration. Thus, we establish a framework that connects microscopic structure, dynamics, local mechanical response, and flow behaviors for supercooled liquids. Finally, we rationalize our framework by leveraging the connection among structure, dynamics, and potential energy landscape (PEL). The PEL model illustrates how local structure, convection and thermal activation collectively determine $\tau_{\mathrm{LC}}$. Notably, it predicts two distinct response groups, which well correspond to the microscopic solid-liquid duality.
@@ I Introduction
For simple liquids [10], Maxwell’s idea has also been introduced to describe the microscopic dynamics [11, 12, 13]. For instance, by constructing a frequency-dependent viscosity with $\tau_{\mathrm{M}}$, one obtains the viscoelastic approximation for transverse-current correlation function [10, 12]. It works well at the molecular level for hard-sphere fluids with moderate densities, while for supercooled states where the viscoelasticity becomes significant, this approximation breaks down. A pronounced challenge for applying Maxwell’s picture to the supercooled state is the dynamic heterogeneity (DH) [14, 15, 16, 17, 18], which implies that different regions in a supercooled liquid exhibit different mobilities. Such spatially heterogeneous mobility renders the global relaxation time ineffective at the local scale [19, 20]. To tackle this problem, one may assign a spatially varying relaxation time [13], whereas the way to extract such local relaxation time can be tricky: It is found that definitions of local time from different considerations may lead to different conclusions even at qualitative level [21, 22]. This problem becomes more complicated when the external flow is applied. In this case, the relaxation of local configuration is a result of the combined effect of thermal activation and convection [23]. Therefore, a proper definition of local characteristic time is expected to distinguish between these two effects.
@@
In this paper, we provide an understanding of the shear rheology of supercooled liquids by embodying Maxwell’s picture of solid-liquid duality at the particle level. With non-equilibrium molecular dynamics (MD) simulations, we introduce the concept of local configurational relaxation time $\tau_{\mathrm{LC}}$, which can be measured once a microscopic configuration is given, regardless of whether the system is in an equilibrium state or under shear. $\tau_{\mathrm{LC}}$ varies in space in the same way as the commonly-understood DH. The competition between the local mobility represented by $\tau_{\mathrm{LC}}$ and the external strain rate determines how the local structure responds to shear, namely, solid-like or liquid-like. In this way, $\tau_{\mathrm{LC}}$ plays a role similar to that of $\tau_{\mathrm{M}}$ in the Maxwell model. By applying this space-dependent solid-liquid duality, we are able to quantitatively understand the rheological response of supercooled liquids to both steady shear and start-up shear in a broad range of shear rates. We introduce $\tau_{\mathrm{LC}}$ in Sec. II, and discuss the flow behaviors of supercooled liquids based on $\tau_{\mathrm{LC}}$ in Sec. III.
Furthermore, we reveal a robust correlation between $\tau_{\mathrm{LC}}$ and microscopic structural features. Our findings suggest a firm relation between relevant local structural parameters and $\tau_{\mathrm{LC}}$, irrespective of shear rates. Based on this correlation, we define a structural order parameter, which effectively distinguishes whether the local response to shear is liquid-like or solid-like. Combining these efforts and the results in Secs. II and III, we bridge the gap between microscopic structure and macroscopic rheology—a long-standing objective in soft condensed matter physics [2, 6, 38]. These results are given in Sec. IV.
In addition, we attempt to formulate the connection from microscopic structure to flow behaviors established in Secs. IIIV by employing the concept of potential energy landscape (PEL) [47, 48, 49, 50]. Within the PEL framework, we illustrate the physical significance of $\tau_{\mathrm{LC}}$ and how it is determined by the collective effect of local structure, convection, and thermal activation. Moreover, we elucidate the mechanism of the solid-liquid duality in response and its consequence in the emergence of nonlinear flow behaviors. The relevant discussion is given in Sec. V.
@@ II.1 Simulation method
where $V$ is the system volume, $N$ is the particle number, $r_{i,x}$ is the $x$ component of the position vector of particle $i$, and $f_{i,y}$ is the $y$ component of the force exerted on particle $i$. Figure 1 shows the viscosities of both samples under steady shear with various shear rates. At large enough $\dot{\gamma}$, both samples exhibit significant shear thinning, which can be expressed by $\eta\sim\dot{\gamma}^{-\lambda}$ [38]. The values of $\lambda$ are, respectively, $0.67$ and $0.66$ for sample A and sample B. For the sake of convenience, we will call the regime featured by $\eta\sim\dot{\gamma}^{-\lambda}$ as the shear-thinning regime, and call the regime between Newtonian regime and the shear-thinning regime as the crossover regime. These rheological regimes are denoted in Fig. 1.
@@ II.2 $\tau_{\mathrm{LC}}$ in quiescent state and local response to start-up shear
Figure 2: (a) Normalized spectra of $\tau_{\mathrm{LC}}$ ($H(\tau_{\mathrm{LC}})$) and $\tau_{\mathrm{sh}}$ ($H(\tau_{\mathrm{sh}})$) of sample A. (b) Normalized spectra of $\tau_{\mathrm{LC}}$ ($H(\tau_{\mathrm{LC}})$) and $\tau_{\mathrm{sh}}$ ($H(\tau_{\mathrm{sh}})$) of sample B. In both (a) and (b), $\tau_{\mathrm{LC}}$ is measured at equilibrium, while $\tau_{\mathrm{sh}}$ is measured under the start-up shear condition with $\dot{\gamma}=0.00025$ for sample A and $\dot{\gamma}=0.00011$ for sample B. (c) Snapshot of the spatial distribution of $\tau_{\mathrm{LC}}$ of sample B at equilibrium. (d) and (e) illustrate the hierarchical feature of the spatial distribution of $\tau_{\mathrm{LC}}$. (d) gives a slice of a fast cluster. It is seen that particles gradually become slower as the distance from the fast center grows. (e) gives a slice of a slow cluster. It is seen that particles gradually become faster as the distance from the slow center grows. Black arrows in (d) and (e) are to guide eyes, indicating that fast particles tend to “grow” from a fast center (d), and slow particles tend to “grow” from a slow center (e).
@@
In supercooled liquids, a particle is strongly restricted by the cage formed by its neighboring particles [56, 57]. Consequently, a particle vibrates inside its cage for most of the time, and occasionally undergoes a cage jump [58]. Knowing particle trajectories, cage jump events can be identified by the algorithm proposed by Candelier et al. [59, 60]. With this method, one can find the time before the first jump from a given time origin for a reference particle $\tau_{\mathrm{p}}$. Here, the subscript “p” denotes “persistence”, in the sense that this definition is consistent with the persistence time in continuous time random walk framework [61, 62]. To enhance the statistics, we employ the isoconfigurational ensemble (ICE) [63], which consists of multiple trajectories from the same configuration but with random initial momenta sampled from the Maxwell-Boltzmann distribution. Note that, by averaging across ICE, one suppresses the uncertainty related to thermal fluctuations. Thus, the ICE-averaged result $\left\langle\tau_{\mathrm{p}}\right\rangle_{\mathrm{ICE}}$ is mainly determined by the configuration at the given time origin, and is irrelevant to the instantaneous momenta. For this reason, we name it local configurational relaxation time, and denote it as $\tau_{\mathrm{LC}}$. Figure 2 (a) and (b) display the normalized equilibrium spectra of $\tau_{\mathrm{LC}}$ of sample A and sample B, respectively. The value of $\tau_{\mathrm{LC}}$ varies from particle to particle, and exhibits significant spatial heterogeneity, as shown in Fig. 2 (c) for sample B at the equilibrium state.
@@
In principle, $\tau_{\mathrm{LC}}$, which represents the local relaxation rate, is relevant to the local response to external load. To explore the relation between $\tau_{\mathrm{LC}}$ and local response, we apply the start-up shear to the equilibrium state, and record the first-jump time from the starting time of this shear for each particle $\tau_{\mathrm{p,sh}}$, where the subscript “sh” represents “shear”. Then, we still perform the ICE average on $\tau_{\mathrm{p,sh}}$ with respect to the configuration at the start of shear, and denote the result as $\tau_{\mathrm{sh}}$ ($\tau_{\mathrm{sh}}=\left\langle\tau_{\mathrm{p,sh}}\right\rangle_{\mathrm{ICE}}$). In Fig. 2 (a) and (b), we plot the distribution of $\tau_{\mathrm{sh}}$ at a given $\dot{\gamma}$ for sample A and sample B, respectively. Compared with $\tau_{\mathrm{LC}}$, the distribution of $\tau_{\mathrm{sh}}$ shifts to lower values, manifesting the facilitation of dynamics by the external shear. For a specific particle, the difference between $\tau_{\mathrm{sh}}$ and $\tau_{\mathrm{LC}}$ reflects the impact of shear on its mobility. Intuitively, slow particles, i.e., particles with large $\tau_{\mathrm{LC}}$, are more susceptible to imposed shear. On the other hand, particles with very small $\tau_{\mathrm{LC}}$ cannot perceive the shear effect, because their spontaneous relaxations are faster than the shear rate.
@@
To quantitatively evaluate the difference between $\tau_{\mathrm{sh}}$ and $\tau_{\mathrm{LC}}$, we perform the following two-step operation for each sample: First, we divide all particles into many bins according to the ascending sequence of $\tau_{\mathrm{LC}}$. In each bin, particles are approximately equal in the value of $\tau_{\mathrm{LC}}$ with a difference up to the bin width. Subsequently, for each bin, we calculate the averaged $\tau_{\mathrm{sh}}$ for particles within that bin. With this operation, we can compare the bin-averaged values of $\tau_{\mathrm{sh}}$ and $\tau_{\mathrm{LC}}$. The results, shown in Fig. 3 (a) and (b) for both samples and various $\dot{\gamma}$, can be summarized as follows: (1) For small $\dot{\gamma}$, $\tau_{\mathrm{sh}}$ is approximately equal to $\tau_{\mathrm{LC}}$ for the majority of $\tau_{\mathrm{LC}}$ values, as indicated by the close alignment with the black thick line denoting $\tau_{\mathrm{sh}}=\tau_{\mathrm{LC}}$. (2) As $\dot{\gamma}$ increases, deviation from this alignment grows at large values of $\tau_{\mathrm{LC}}$, suggesting that more and more particles are accelerated by shear. This deviation is pronounced in the shear-thinning regime, manifested by the flattening of the $\tau_{\mathrm{sh}}(\tau_{\mathrm{LC}})$ curve. These results are consistent with our intuition.
@@
where $\tau_{2}$ is the non-Gaussian time at equilibrium, which measures the waiting time for the system to exhibit significant DH [64]. According to the preceding criterion, slow particles exhibit distinct relaxation heterogeneity compared with the equilibrium situation, thereby being considered as being strongly influenced by shear. Then, the fast particles are identified by satisfying $\tau_{\mathrm{LC}}-\tau_{\mathrm{sh}}<\tau_{2}$. The boundary between fast and slow groups $\tau_{\mathrm{LC,c}}$ ($\tau_{\mathrm{LC,c}}=\tau_{\mathrm{sh}}+\tau_{2}$) is denoted by red vertical bar for each $\dot{\gamma}$ in Fig. 3 (a) and (b), where the length of the bar is set to $\tau_{2}$. $\tau_{\mathrm{LC,c}}$ reflects the time scale of external shear, so that it is expected to linearly depend on $\dot{\gamma}^{-1}$ [65]. This linearity is shown in Fig. 3 (c) and (d), implying that our way of determining $\tau_{\mathrm{LC,c}}$ is reasonable.
@@
Figure 4: (a) Slice of the spatial distribution of $\tau_{\mathrm{LC}}$ of sample A at equilibrium. (b) – (d) Slices of the spatial distribution of $\tau_{\mathrm{sh}}$ of sample A obtained under start-up shear with various $\dot{\gamma}$ denoted at each panel. In (b) – (d), white dashed lines denote the boundaries between fast and slow groups of particles found by Eq. 2. (e) – (h) show the results for sample B.
@@
The spatial distribution of fast or slow particles and its variation with $\dot{\gamma}$ are of particular importance for clarifying the local response. To understand this point, we first notice that the distribution of $\tau_{\mathrm{LC}}$ is hierarchical. Taking fast particles as an example: Fast particles tend to aggregate, as a manifestation of DH. The center of an aggregation is formed by a few very fast particles. As the distance from the center increases, the particles gradually become slower, as shown in Fig. 2 (d). This is consistent with the dynamic facilitation mechanism, which suggests that local relaxation events propagate to neighboring regions [66, 59, 60, 67]. Similarly, slow particles exhibit clustering with centers formed by a few very slow particles. Within a slow cluster, particles gradually become faster as the distance from center increases, as shown in Fig. 2 (e). Analogous pattern is also observed in a recent study [68]. When a start-up shear is applied, the local response is reflected by the distribution of $\tau_{\mathrm{sh}}$. According to the previous analysis, as $\dot{\gamma}$ increases, more particles will be involved as slow ones, manifested by the expansion of slow clusters. This trend is clearly seen in Fig. 4. Figure 4 (a) shows the spatial distribution of $\tau_{\mathrm{LC}}$ for sample A at equilibrium. Figure 4 (b) – (d) show the spatial distributions of $\tau_{\mathrm{sh}}$ of sample A under start-up shear with various $\dot{\gamma}$. The boundaries of slow clusters, found by $\tau_{\mathrm{sh,b}}=\tau_{\mathrm{LC,c}}-\tau_{2}$, are denoted by dashed lines (the detailed method for circling these clusters is given in Appendix A). It is seen that $\tau_{\mathrm{LC}}$ and $\tau_{\mathrm{sh}}$ exhibit similar heterogeneous features. Moreover, the boundaries of slow clusters expand as $\dot{\gamma}$ increases. Same observation is found for sample B, as shown in Fig. 4 (e) – (h).
@@
For the meaning of “slow” and “fast”, we would like to discuss the following two points. (i) As $\dot{\gamma}$ increases, the mean value of $\tau_{\mathrm{sh}}$ decreases, suggesting faster dynamics. On the other hand, more particles are involved in slow regions. These two observations are not contradictory. When we identify a group of particles as slow, it means that their mobility is slower than external shear. Thus, as $\dot{\gamma}$ increases, the dynamics is facilitated by shear, meanwhile, more particles become slower than the external shear and are identified as slow ones. (ii) Considering that the spectra of $\tau_{\mathrm{sh}}$ and $\tau_{\mathrm{LC}}$ are not bimodal, one may feel that the dichotomy used here is too rough. Especially, for particles with mediate mobility, whether they are fast or slow is obscure. This problem can be clarified by the clustering feature of slow/fast regions. As discussed in the preceding paragraph, typically, a slow (or fast) cluster has a center formed by a few very slow (or fast) particles. As the distance from the center increases, particles gradually become faster (or slower). Consequently, the particles with mediate mobility are mainly at the boundaries between fast and slow clusters. Whether they are identified as slow or fast may result in small fluctuation of the boundaries, but will not influence the identification of a cluster to be slow or fast.

### v2

@@ Abstract
Understanding the flow behaviors of supercooled liquids presents a major challenge in liquid-state physics due to the strong nonlinearity and rich phenomena. To unravel this complexity, we introduce the concept of local configurational relaxation time $\tau_{\mathrm{LC}}$, which allows us to embody the solid-liquid duality, proposed by Maxwell for phenomenologically describing materials’ response to external load, at the particle level. The spatial distribution of $\tau_{\mathrm{LC}}$ in flow is heterogeneous. Depending on the comparison between the local mobility measured by $\tau_{\mathrm{LC}}$ and the external shear rate, the shear response of local regions is either solid-like or liquid-like. In this way, $\tau_{\mathrm{LC}}$ plays a role similar to the Maxwell time. By applying this microscopic solid-liquid duality to different conditions of shear flow with a wide range of shear rates, we describe the emergence of shear thinning in steady shear, and predict the major characteristics of the transient response to start-up shear. Furthermore, we reveal a clear structural foundation for $\tau_{\mathrm{LC}}$ and the solid-liquid duality associated with it by introducing an order parameter extracted from local configuration. Thus, we establish a framework that connects microscopic structure, dynamics, local mechanical response, and flow behaviors for supercooled liquids. Finally, we rationalize our framework in terms of activations from energy basins that are facilitated by shear. This model illustrates how local structure, convection and thermal activation collectively determine $\tau_{\mathrm{LC}}$. Notably, it predicts two distinct response groups, which well correspond to the microscopic solid-liquid duality.
@@ I Introduction
For simple liquids [10], Maxwell’s idea has also been introduced to describe the microscopic dynamics [11, 12, 13]. For instance, by constructing a frequency-dependent viscosity with $\tau_{\mathrm{M}}$, one obtains the viscoelastic approximation for transverse-current correlation function [10, 12]. It works well at the molecular level for hard-sphere fluids with moderate densities, while for supercooled states where the viscoelasticity becomes significant, this approximation breaks down. A pronounced challenge for applying Maxwell’s picture to the supercooled state is the dynamic heterogeneity (DH) [14, 15, 16, 17, 18], which implies that different regions in a supercooled liquid exhibit different mobilities. Such spatially heterogeneous mobility renders the global relaxation time ineffective at the local scale [19, 20]. To tackle this problem, one may assign a spatially varying relaxation time [13], whereas the way to extract such local relaxation time can be tricky: It is found that definitions of local time from different considerations may lead to different conclusions even at qualitative level [21, 22]. This problem becomes more complicated when the external flow is applied. In this case, the relaxation of local configuration is a result of the combined effect of thermal activation and convection [23, 24, 25]. Therefore, a proper definition of local characteristic time is expected to distinguish between these two effects.
@@
In this paper, we provide an understanding of the shear rheology of supercooled liquids by embodying Maxwell’s picture of solid-liquid duality at the particle level. With non-equilibrium molecular dynamics (MD) simulations, we introduce the concept of local configurational relaxation time $\tau_{\mathrm{LC}}$, which can be measured once a microscopic configuration is given, regardless of whether the system is in an equilibrium state or under shear. $\tau_{\mathrm{LC}}$ varies in space in the same way as the commonly-understood DH. The competition between the local mobility represented by $\tau_{\mathrm{LC}}$ and the external strain rate determines how the local structure responds to shear, namely, solid-like or liquid-like. In this way, $\tau_{\mathrm{LC}}$ plays a role similar to that of $\tau_{\mathrm{M}}$ in the Maxwell model. By applying this space-dependent solid-liquid duality, we are able to quantitatively understand the rheological response of supercooled liquids to both steady shear and start-up shear in a broad range of shear rates. We introduce $\tau_{\mathrm{LC}}$ in section II, and discuss the flow behaviors of supercooled liquids based on $\tau_{\mathrm{LC}}$ in section III.
Furthermore, we reveal a robust correlation between $\tau_{\mathrm{LC}}$ and microscopic structural features. Our findings suggest a firm relation between relevant local structural parameters and $\tau_{\mathrm{LC}}$, irrespective of shear rates. Based on this correlation, we define a structural order parameter, which effectively distinguishes whether the local response to shear is liquid-like or solid-like. Combining these efforts and the results in sections II and III, we bridge the gap between microscopic structure and macroscopic rheology—a long-standing objective in soft condensed matter physics [2, 6, 40]. These results are given in section IV.
In addition, we formulate the connection from the microscopic structure to flow behaviors established in sections IIIV based on a simple picture of shear-facilitated activation from basins (local configuration for particle) [28, 23, 49]. With this model, we illustrate the physical significance of $\tau_{\mathrm{LC}}$ and how it is determined by the collective effect of local structure, convection, and thermal activation. Moreover, we elucidate the mechanism of the solid-liquid duality in response and its consequence in the emergence of nonlinear flow behaviors. The relevant discussion is given in section V.
@@ II.1 Simulation method
where $V$ is the system volume, $N$ is the particle number, $r_{i,x}$ is the $x$ component of the position vector of particle $i$, and $f_{i,y}$ is the $y$ component of the force exerted on particle $i$. Figure 1 shows the viscosities of both samples under steady shear with various shear rates. At large enough $\dot{\gamma}$, both samples exhibit significant shear thinning, which can be expressed by $\eta\sim\dot{\gamma}^{-\lambda}$ [40]. The values of $\lambda$ are, respectively, $0.67$ and $0.66$ for sample A and sample B. For the sake of convenience, we will call the regime featured by $\eta\sim\dot{\gamma}^{-\lambda}$ as the shear-thinning regime, and call the regime between Newtonian regime and the shear-thinning regime as the crossover regime. These rheological regimes are denoted in figure 1.
@@ II.2 $\tau_{\mathrm{LC}}$ in quiescent state and local response to start-up shear
Figure 2: (a) Normalized distribution of $\tau_{\mathrm{LC}}$ ($H(\tau_{\mathrm{LC}})$) and $\tau_{\mathrm{sh}}$ ($H(\tau_{\mathrm{sh}})$) of sample A. (b) Normalized distribution of $\tau_{\mathrm{LC}}$ ($H(\tau_{\mathrm{LC}})$) and $\tau_{\mathrm{sh}}$ ($H(\tau_{\mathrm{sh}})$) of sample B. In both (a) and (b), $\tau_{\mathrm{LC}}$ is measured at equilibrium, while $\tau_{\mathrm{sh}}$ is measured under the start-up shear condition with $\dot{\gamma}=0.00025$ for sample A and $\dot{\gamma}=0.00011$ for sample B. (c) Snapshot of the spatial distribution of $\tau_{\mathrm{LC}}$ of sample B at equilibrium. (d) and (e) illustrate the hierarchical feature of the spatial distribution of $\tau_{\mathrm{LC}}$. (d) gives a slice of a fast cluster. It is seen that particles gradually become slower as the distance from the fast center grows. (e) gives a slice of a slow cluster. It is seen that particles gradually become faster as the distance from the slow center grows. Black arrows in (d) and (e) are to guide eyes, indicating that fast particles tend to “grow” from a fast center (d), and slow particles tend to “grow” from a slow center (e).
@@
In supercooled liquids, a particle is strongly restricted by the cage formed by its neighboring particles [55, 56]. Consequently, a particle vibrates inside its cage for most of the time, and occasionally undergoes a cage jump [57, 58]. Knowing particle trajectories, cage jump events can be identified by the algorithm proposed by Candelier et al. [59, 60] (Appendix A.1). With this method, one can find the time before the first jump from a given time origin for a reference particle $\tau_{\mathrm{p}}$. Here, the subscript “p” denotes “persistence”, in the sense that this definition is consistent with the persistence time in continuous time random walk framework [61, 62]. To enhance the statistics, we employ the isoconfigurational ensemble (ICE) [63], which consists of multiple trajectories from the same configuration but with random initial momenta sampled from the Maxwell-Boltzmann distribution. Note that, by averaging across ICE, one suppresses the uncertainty related to thermal fluctuations. Thus, the ICE-averaged result $\left\langle\tau_{\mathrm{p}}\right\rangle_{\mathrm{ICE}}$ is mainly determined by the configuration at the given time origin, and is irrelevant to the instantaneous momenta. For this reason, we name it local configurational relaxation time, and denote it as $\tau_{\mathrm{LC}}$. Figure 2 (a) and (b) display the normalized equilibrium distribution of $\tau_{\mathrm{LC}}$ of sample A and sample B, respectively. The value of $\tau_{\mathrm{LC}}$ varies from particle to particle, and exhibits significant spatial heterogeneity, as shown in figure 2 (c) for sample B at the equilibrium state.
@@
In principle, $\tau_{\mathrm{LC}}$, which represents the local relaxation rate, is relevant to the local response to external load. To explore the relation between $\tau_{\mathrm{LC}}$ and local response, we apply the start-up shear to the equilibrium state, and record the first-jump time from the starting time of this shear for each particle $\tau_{\mathrm{p,sh}}$, where the subscript “sh” represents “shear”. Then, we still perform the ICE average on $\tau_{\mathrm{p,sh}}$ with respect to the configuration at the start of shear, and denote the result as $\tau_{\mathrm{sh}}$ ($\tau_{\mathrm{sh}}=\left\langle\tau_{\mathrm{p,sh}}\right\rangle_{\mathrm{ICE}}$). In figure 2 (a) and (b), we plot the distribution of $\tau_{\mathrm{sh}}$ at a given $\dot{\gamma}$ for sample A and sample B, respectively. Compared with $\tau_{\mathrm{LC}}$, the distribution of $\tau_{\mathrm{sh}}$ shifts to lower values, manifesting the facilitation of dynamics by the external shear. For a specific particle, the difference between $\tau_{\mathrm{sh}}$ and $\tau_{\mathrm{LC}}$ reflects the impact of shear on its mobility. Intuitively, slow particles, i.e., particles with large $\tau_{\mathrm{LC}}$, are more susceptible to imposed shear. On the other hand, particles with very small $\tau_{\mathrm{LC}}$ cannot perceive the shear effect, because their spontaneous relaxations are faster than the shear rate.
@@
To quantitatively evaluate the difference between $\tau_{\mathrm{sh}}$ and $\tau_{\mathrm{LC}}$, we perform the following two-step operation for each sample: First, we divide all particles into many bins according to the ascending sequence of $\tau_{\mathrm{LC}}$. In each bin, particles are approximately equal in the value of $\tau_{\mathrm{LC}}$ with a difference up to the bin width. Subsequently, for each bin, we calculate the averaged $\tau_{\mathrm{sh}}$ for particles within that bin. With this operation, we can compare the bin-averaged values of $\tau_{\mathrm{sh}}$ and $\tau_{\mathrm{LC}}$. The results, shown in figure 3 (a) and (b) for both samples and various $\dot{\gamma}$, can be summarized as follows: (1) For small $\dot{\gamma}$, $\tau_{\mathrm{sh}}$ is approximately equal to $\tau_{\mathrm{LC}}$ for the majority of $\tau_{\mathrm{LC}}$ values, as indicated by the close alignment with the black thick line denoting $\tau_{\mathrm{sh}}=\tau_{\mathrm{LC}}$. (2) As $\dot{\gamma}$ increases, deviation from this alignment grows at large values of $\tau_{\mathrm{LC}}$, suggesting that more and more particles are accelerated by shear. This deviation is pronounced in the shear-thinning regime, manifested by the flattening of the $\tau_{\mathrm{sh}}(\tau_{\mathrm{LC}})$ curve. These results are consistent with our intuition.
@@
where $\tau_{2}$ is the non-Gaussian time at equilibrium, which measures the waiting time for the system to exhibit significant DH [64]. According to the preceding criterion, slow particles exhibit distinct relaxation heterogeneity compared with the equilibrium situation, thereby being considered as being strongly influenced by shear. Then, the fast particles are identified by satisfying $\tau_{\mathrm{LC}}-\tau_{\mathrm{sh}}<\tau_{2}$. The boundary between fast and slow groups $\tau_{\mathrm{LC,c}}$ ($\tau_{\mathrm{LC,c}}=\tau_{\mathrm{sh}}+\tau_{2}$) is denoted by red vertical bar for each $\dot{\gamma}$ in figure 3 (a) and (b), where the length of the bar is set to $\tau_{2}$. $\tau_{\mathrm{LC,c}}$ reflects the time scale of external shear, so that it is expected to linearly depend on $\dot{\gamma}^{-1}$ [65]. This linearity is shown in figure 3 (c) and (d), implying that our way of determining $\tau_{\mathrm{LC,c}}$ is reasonable.
@@
Figure 4: (a) Slice of the spatial distribution of $\tau_{\mathrm{LC}}$ of sample A at equilibrium. (b) – (d) Slices of the spatial distribution of $\tau_{\mathrm{sh}}$ of sample A obtained under start-up shear with various $\dot{\gamma}$ denoted at each panel. In (b) – (d), white dashed lines denote the boundaries between fast and slow groups of particles found by equation 2. (e) – (h) show the results for sample B.
@@
The spatial distribution of fast or slow particles and its variation with $\dot{\gamma}$ are of particular importance for clarifying the local response. To understand this point, we first notice that the distribution of $\tau_{\mathrm{LC}}$ is hierarchical. Taking fast particles as an example: Fast particles tend to aggregate, as a manifestation of DH. The center of an aggregation is formed by a few very fast particles. As the distance from the center increases, the particles gradually become slower, as shown in figure 2 (d). This is consistent with the dynamic facilitation mechanism, which suggests that local relaxation events propagate to neighboring regions [66, 59, 60, 67]. Similarly, slow particles exhibit clustering with centers formed by a few very slow particles. Within a slow cluster, particles gradually become faster as the distance from center increases, as shown in figure 2 (e). Analogous pattern is also observed in a recent study [68]. When a start-up shear is applied, the local response is reflected by the distribution of $\tau_{\mathrm{sh}}$. According to the previous analysis, as $\dot{\gamma}$ increases, more particles will be involved as slow ones, manifested by the expansion of slow clusters. This trend is clearly seen in figure 4. Figure 4 (a) shows the spatial distribution of $\tau_{\mathrm{LC}}$ for sample A at equilibrium. Figure 4 (b) – (d) show the spatial distributions of $\tau_{\mathrm{sh}}$ of sample A under start-up shear with various $\dot{\gamma}$. The boundaries of slow clusters, found by $\tau_{\mathrm{sh,b}}=\tau_{\mathrm{LC,c}}-\tau_{2}$, are denoted by dashed lines (the detailed method for circling these clusters is given in Appendix A). It is seen that $\tau_{\mathrm{LC}}$ and $\tau_{\mathrm{sh}}$ exhibit similar heterogeneous features. Moreover, the boundaries of slow clusters expand as $\dot{\gamma}$ increases. Same observation is found for sample B, as shown in figure 4 (e) – (h).
@@
For the meaning of “slow” and “fast”, we would like to discuss the following two points. (i) As $\dot{\gamma}$ increases, the mean value of $\tau_{\mathrm{sh}}$ decreases, suggesting faster dynamics. On the other hand, more particles are involved in slow regions. These two observations are not contradictory. When we identify a group of particles as slow, it means that their mobility is slower than external shear. Thus, as $\dot{\gamma}$ increases, the dynamics is facilitated by shear, meanwhile, more particles become slower than the external shear and are identified as slow ones. (ii) Considering that the distributions of $\tau_{\mathrm{sh}}$ and $\tau_{\mathrm{LC}}$ are not bimodal, one may feel that the dichotomy used here is too rough. Especially, for particles with mediate mobility, whether they are fast or slow is obscure. This problem can be clarified by the clustering feature of slow/fast regions. As discussed in the preceding paragraph, typically, a slow (or fast) cluster has a center formed by a few very slow (or fast) particles. As the distance from the center increases, particles gradually become faster (or slower). Consequently, the particles with mediate mobility are mainly at the boundaries between fast and slow clusters. Whether they are identified as slow or fast may result in small fluctuation of the boundaries, but will not influence the identification of a cluster to be slow or fast.